Please help me solve this equation:
1/2x+3y=5/6
1/3x-5y=16/9

Answer by rothauserc(4718)

The problem states
Please help me solve this equation:
1/2x+3y=5/6
1/3x-5y=16/9
we solve this using substitution since we have two equations and two unknowns
using equation one we multiply both sides of the = by 2 and we get
x+6y=5/3 and
x=5/3 - 6y
using the second equation we multiply both sides of the = by 3 and we get
x-15y=16/3
now substitute x with 5/3 - 6y and we get
5/3 - 6y - 15y = 16/3
-21y = 11/3
and y = (11/3) / -21 = -.17
and 1/2 * x + 3*-.17 = 5/6
multiply by 2 and we get
x + 6*-.17 = 5/3
x-1.02 = 1.67 and x = 2.69